# Learning to Inference Adaptively for Multimodal Large Language Models

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have shown impressive capabilities in reasoning, yet come with substantial computational cost, limiting their deployment in resource-constrained settings. Despite recent efforts on improving the efficiency of MLLMs, prior solutions fall short in responding to varying runtime conditions, in particular changing resource availability (e.g., contention due to the execution of other programs on the device). To bridge this gap, we introduce AdaLLaVA, an adaptive inference framework that learns to dynamically reconfigure operations in an MLLM during inference, accounting for the input data and a latency budget. We conduct extensive experiments across benchmarks involving question-answering, reasoning, and hallucination. Our results show that AdaLLaVA effectively adheres to input latency budget, achieving varying accuracy and latency tradeoffs at runtime. Further, we demonstrate that AdaLLaVA adapts to both input latency and content, can be integrated with token selection for enhanced efficiency, and generalizes across MLLMs. Our project webpage with code release is at https://zhuoyan-xu.github.io/ada-llava/.
@@ 1 Introduction
Large language models (LLMs) [43, 2] have recently been extended to connect visual and textual data, giving rise to multimodal large language models (MLLMs). Exemplified by LLaVA [35, 36] and other works [37, 30, 32, 66, 1, 54], MLLMs have shown impressive visual reasoning capabilities, but come with significant computational costs. Several efforts have sought to improve the efficiency of MLLMs by exploring lightweight architectures, mixture of experts, or token selection techniques [66, 61, 34, 49, 8]. A common characteristic of these MLLM methods is that they generate models with a fixed accuracy and latency footprint when performing inference on a given input.
@@
We argue that MLLMs with fixed computational footprints are insufficient for real-world deployment. Consider an example of deploying an MLLM on a server farm. Different requests may have different latency requirements, e.g., requests from a mobile application require instant feedback to users, while those from a recommendation system can tolerate higher latency due to less frequent updates. Further, the available computing resources may vary over time as the overall loads of the system fluctuate. Similarly, when deployed on an edge device, the latency budget often remains constant, yet the computing resources may vary due to contention produced by other concurrent programs. In spite of this need, developing adaptive inference strategies for MLLMs that are robust across varying computational budgets [25] remains a key research challenge.
To bridge this gap, we propose latency-aware adaptive inference for MLLMs, aiming to dynamically adjust a model’s computational load based on input content and a specified latency or compute budget [^1] In this paper, we measure a model’s latency and its budget using the number of floating-point operations (FLOPs). Thus, the terms “compute budget” and “latency budget” are used interchangeably throughout.. This problem is of both conceptual interest and practical significance. Our key insight is that a modern MLLM can be viewed as a collection of shallower models, where choosing among these models enables dynamic reconfiguration during inference. For example, prior works have shown that Transformer blocks in an LLM and some attention heads within these blocks can be bypassed with minimal impact on accuracy, while reducing latency [52, 5, 26]. Therefore, strategically selecting these operations during inference results in a set of models with shared parameters but distinct accuracy-latency tradeoffs, allowing the MLLM to flexibly respond to varying latency budgets.
@@
We conduct extensive experiments to evaluate AdaLLaVA. Our results demonstrate that AdaLLaVA can achieve a range of accuracy-latency tradeoffs at runtime. AdaLLaVA exhibits strong adaptability to different latency budgets, effectively trading accuracy for compute during inference. Across several benchmarks, AdaLLaVA retains comparable performance to its base MLLM while operating with higher efficiency (see Fig. 1). For example, on several comprehensive benchmarks, AdaLLaVA can achieve 99.0% and 98.2% average performance of the baseline model when using only 80% and 65% of the latency budget, respectively. Importantly, it consistently adheres to specified latency constraints and generates content-aware execution plans tailored to input images. Furthermore, we show that AdaLLaVA can be integrated with token selection techniques designed to enhance efficiency, making it a versatile solution for adaptive inference in MLLMs.
Our key contributions are summarized into three folds.
@@ 2 Related Work
Adaptive inference. Adaptive inference refers to the capability in which the computational complexity of making predictions is dynamically adjusted based on the input data, latency budget, or desired accuracy levels [18]. Early works focused on the selection of hand-crafted features in multi-stage prediction pipelines [25, 60, 15]. More recent works have extended these ideas to deep models. For convolutional networks, methods have been developed to selectively downsample inputs, skip layers, or exist early during inference [12, 31, 55, 4, 59, 19, 24, 42]. For vision Transformers, various approaches have been proposed, such as selecting different image patches [56, 46, 44], and choosing different attention heads and blocks [41, 26]. Similar ideas have also been explored for LLMs, where models selectively process tokens [47] or execute a subset of the operations [11, 48] during inference.
Our approach is conceptually similar to exisiting methods by dynamically selecting a subset of model components during inference. Yet unlike prior methods, our work specifically targets the latency-aware inference of MLLMs, predicting feasible execution plans tailored for input while adhering to varying latency budgets.
@@
While our approach also aims to improve the efficiency of MLLMs, it focuses on dynamically adjusting an MLLM to fit varying latency budgets during inference. This makes our approach orthogonal to prior efforts on developing inherently more efficient MLLMs. Through our experiments, we will demonstrate that our approach is compatible with lightweight models and integrates seamlessly with existing token-pruning techniques (e.g.[49, 8]).
@@ 3.1 Preliminaries: MLLMs
Figure 2: Overview of AdaLLaVA. (a) Model architecture: Our latency encoder embeds an input latency budget into a latency token, which is further processed by the early part of the LLM. The resulting embedding is then fed into the scheduler, leading to the output of an execution plan that control individual operations in the remaining part of the LLM. Our latency encoder and scheduler are jointly trained with the MLLM. (b) AdaLLaVA-L: the scheduler controls the execution of entire Transformer blocks. (c) AdaLLaVA-H: the scheduler controls the execution of attention heads and MLP neurons, by masking out their activation values and the corresponding weights.
@@ 3.3 Learning to Schedule Execution Plans
Deterministic modeling. A possible solution is to solve a relaxed version of the constrained optimization at training time. We initially explored this solution, where we task $g(\cdot)$ to predict a hard execution plan with binary switches $\mathbf{s}$ and attribute latency violation as part of the objective. This leads to the following loss
@@ 3.5 Model Instantiation
Model architecture. We aim to keep the scheduler as light as possible to minimize the FLOPs overhead while ensuring that it remains expressive enough to make optimal decisions. To this end, we reuse part of the LLM $f(\cdot)$ to extract visual-language features and encode the latency constraint for the scheduler. Specifically, we first design a latency encoder that converts a latency budget into a token embedding, which is then appended to the original input sequence before being processed by the LLM layers. Within the LLM, the latency token is processed by a few Transformer blocks, attending to all visual-language tokens. The processed token is then passed to a lightweight scheduler that generates the execution plan for the rest of the LLM. Notably, the first few Transformer blocks in the LLM serve two purposes: it simultaneously processes regular MLLM tasks and learns resource allocation based on both content and budget constraints. This design is depicted in Fig. 2 (a).
@@
BLIP-2 [32] Vicuna-13B - - - 41.0 61 42.5 85.3 1293.8 -
InstructBLIP [10] Vicuna-7B - - - - 60.5 50.1 - - 36.0
InstructBLIP [10] Vicuna-13B - - - - 63.1 50.7 78.9 1212.8 -
Shikra [7] Vicuna-13B - - - 77.4 - - - - 58.8
IDEFICS-9B [28] LLaMA-7B - - - 50.9 - 25.9 - - 48.2
IDEFICS-80B [28] LLaMA-65B - - - 60.0 - 30.9 - - 54.5
Qwen-VL [3] Qwen-7B - - - 78.8 67.1 63.8 - - 38.2
Qwen-VL-Chat [3] Qwen-7B - - - 78.2 68.2 61.5 - 1487.5 60.6
@@ 4.1 Experimental Setup
Experiment protocol. In most of our experiments, we build on LLaVA-1.5 [36]. Training LLaVA and many other MLLMs typically involves two stages: (1) vision-language alignment pre-training; and (2) visual instruction tuning. We focus on the second stage and seek to jointly finetune the LLM within the MLLM and train our scheduler using visual instruction data, while keeping the vision encoder frozen. Once trained, we perform zero-shot inference across multiple benchmarks following the common practice in the community [36], but under varying latency budgets.
@@
Baselines and model variants. Baselines in our main results include BLIP2 [32, 10], Shikra [7], IDEFICS [28], Qwen-VL [3], and LLaVA-1.5 [36]. Our closest competitor is the base model LLaVA-1.5 [36]. We evaluate AdaLLaVA with 7B and 13B (see supplement) models, and across two different designs: (a) AdaLLaVA-L for selecting Transformer blocks; and (b) AdaLLaVA-H for selecting attention heads and MLP activations. In our additional analyses, we also use Mipha-3B [67] as the base model.
@@ 4.2 Main Results
Importantly, for all results, AdaLLaVA adheres to the specified latency budgets, as we will discuss in Sec. 4.3.
Integration with token selection. Token selection techniques have demonstrated recent success in improving the efficiency of MLLMs [49, 8]. AdaLLaVA presents an orthogonal direction in adaptive inference. We now demonstrate that AdaLLaVA can be integrated with token selection to further enhance the efficiency. We combine AdaLLaVA with PruMerge+ [49] and FastV [8], two latest token selection methods designed for MLLMs. For FastV, we set filtering layer $K$=2 and filtering ratio $R$=50to ensure consistent comparison. The results are shown in Tab. 1.
When integrated with PruMerge+ and FastV, AdaLLaVA shows significantly improved efficiency across board, when compared to AdaLLaVA with LLaVA-1.5. Again, AdaLLaVA adapts to varying latency budgets and achieves competitive performance relative to the base model (PruMerge+/FastV). For example, with PruMerge+, AdaLLaVA-H shows 2.45% average performance boost at 85% compute budget and only 1.01% performance drop at 60%. A surprising observation is that AdaLLaVA-H achieves strong performance at 85% latency budget, sometimes beating the base model with token pruning. Overall, our results suggest that AdaLLaVA complements to existing token selection approaches. When integrated with Prumerge+ at an 85% latency budget, our approach reduces computational requirements by 70% while maintaining performance with only 1.7% drop in accuracy.
@@ 4.3 Additional Analyses
Our results shows that AdaLLaVA can empower a base MLLM with static compute footprint (i.e., LLaVA-1.5, PruMerge+, or FastV as individual dots in the Fig. 3) to adapt to varying accuracy-latency tradeoffs (i.e., the corresponding curves in Fig. 3). With varying latency budgets from 50% to 100%, AdaLLaVA effectively trades compute with accuracy. Integrating with token selection methods (PruMerge+ / FastV) further improves the overall efficiency. Thanks to our sampling process in the probabilistic modeling, AdaLLaVA maintains 0% latency violation.
@@
We visualize attention maps from the latency token to all input visual tokens, computed right before the latency token is fed into the scheduler. This is shown in Fig. 4. These attention maps point to key regions in the input image that are necessary for answering the target question. For example, in the top, attention concentrates on “Yes Man” for the movie title question but shifts to the actor name for actor identification question. Further, we visualize the execution plans of different input content given by our scheduler in Fig. 5. Under the same latency budget, AdaLLaVA generates distinct execution plans conditioned on the input visual content. These results show AdaLLaVA’s ability to dynamically adjust its computational focus and execution plan based on the input image and text query. See our supplement for additional visualizations.
Generalization across MLLMs. We further demonstrate that AdaLLaVA can generalize to other MLLMs beyond LLaVA. We consider Mipha-3B [67], a lightweight MLLM built on Phi-2.7B [23]. Specifically, we apply AdaLLaVA-L on Mipha-3B, following its training strategy [67], and report the results on MME benchmark, shown in Tab. 2. These results have similar trend to those with LLaVA-1.5 in Tab. 1.
@@
Additional ablations. Ablations on (1) the number and granularity of switches; and (2) different designs of switches (i.e., AdaLLaVA-H vs. AdaLLaVA-L) are included in our supplement due to the space limit.
@@ 5 Conclusion and Discussion
Adaptive inference of MLLMs. Unlike LLMs, MLLMs include a vision encoder and process a large number of redundant visual tokens. While our paper focus on the scheduling of the LLM component, this adaptivity can be further extended to token selection and vision encoder. We hope this work will provide a step toward making MLLMs more viable for real-world applications where computational resources may fluctuate significantly.
Relationship to other efficiency methods. This paper explores adaptive inference in MLLMs, emphasizing adaptability to varying latency budgets in a single model. Our approach is orthogonal to prior methods aimed at improving inference efficiency, such as sparse attention [9] and token selection [49]. Indeed, many of these techniques (e.g. token selection as shown in the paper) can be integrated with our framework to further enhance efficiency.
Practical deployment. Our work focuses on algorithm-level innovation, leaving system-level serving optimization as future work. Conceptually, serving AdaLLaVA is similar to serving MoE-based LLMs [20], which also dynamically route tokens to different execution paths based on the input. We invite joint effort from the vision, learning, and systems communities to further explore this direction.
@@
In this supplementary material, we provide (1) additional implementation details (see Supp. A); (2) detailed results, including broader benchmarks and other LLM backbones on AdaLLaVA accompanying our experiments in Sec. 4 (see Supp. B); (3) further ablations on design of switches (see Supp. C) and (4) additional qualitative results on latency and content adaptivity (see Supp. D) We hope that this document will complement our main paper.

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) have shown impressive capabilities in visual reasoning, yet come with substantial computational cost, limiting their deployment in resource-constrained settings. Despite recent effort on improving the efficiency of MLLMs, prior solutions fall short in responding to varying runtime conditions, in particular changing resource availability (e.g., contention due to the execution of other programs on the device). To bridge this gap, we introduce AdaLLaVA, an adaptive inference framework that learns to dynamically reconfigure operations in an MLLM during inference, accounting for the input data and a latency budget. We conduct extensive experiments across benchmarks involving question-answering, reasoning, and hallucination. Our results show that AdaLLaVA effectively adheres to input latency budget, achieving varying accuracy and latency tradeoffs at runtime. Further, we demonstrate that AdaLLaVA adapts to both input latency and content, can be integrated with token selection for enhanced efficiency, and generalizes across MLLMs. Our project webpage with code release is at https://zhuoyan-xu.github.io/ada-llava/.
@@ 1 Introduction
Large language models (LLMs) [43, 2] have recently been extended to connect visual and textual data, giving rise to multimodal large language models (MLLMs). Exemplified by LLaVA [35, 36] and other works [37, 30, 32, 69, 1, 54], MLLMs have shown impressive visual reasoning capabilities, but come with significant computational costs. Several efforts have sought to improve the efficiency of MLLMs by exploring lightweight architectures, mixture of experts, or token selection techniques [69, 63, 34, 49, 8]. However, prior approaches typically exhibit a fixed accuracy and latency footprint during inference, rendering them incapable of adapting to varying compute budget or input content.
@@
We argue that MLLMs with fixed computational footprints are insufficient for real-world deployment. Consider an example of deploying an MLLM on a server farm. Different requests may have different latency requirements, e.g., requests from a mobile application require instant feedback to users, while asynchronous processing tasks such as video summarization can tolerate higher latency due to their non-interactive nature. Further, the available computing resources may vary over time as the overall load on the system fluctuates. Similarly, when deployed on an edge device, the latency budget often remains constant, yet the computing resources may vary due to contention produced by other concurrent programs. In spite of this need, developing adaptive inference strategies for MLLMs that are robust across varying computational budgets [25] remains an open research challenge.
To bridge this gap, we propose latency-aware adaptive inference for MLLMs, aiming to dynamically adjust a model’s computational load based on input content and a specified latency or compute budget. [^1] In this paper, we measure a model’s latency and its budget using the number of floating-point operations (FLOPs). Thus, the terms “compute budget” and “latency budget” are used interchangeably throughout. This problem is of both conceptual interest and practical significance. Our key insight is that a modern MLLM can be viewed as a collection of shallower models, where choosing among these models enables dynamic reconfiguration during inference. For example, prior works have shown that Transformer blocks in an LLM and some attention heads within these blocks can be bypassed with minimal impact on accuracy, while reducing latency [52, 5, 26]. Therefore, strategically selecting these operations during inference results in a set of models with shared parameters but distinct accuracy-latency tradeoffs, allowing the MLLM to flexibly respond to varying latency budgets and content complexity.
@@
We conduct extensive experiments to evaluate AdaLLaVA. Our results demonstrate that AdaLLaVA can achieve a range of accuracy-latency tradeoffs at runtime. AdaLLaVA exhibits strong adaptability to different latency budgets, effectively trading accuracy for compute during inference. Across several benchmarks, AdaLLaVA retains comparable performance to its base MLLM while operating with higher efficiency (see Fig. 1). For example, on several comprehensive benchmarks, AdaLLaVA can achieve 99.0% and 98.2% average performance of the baseline LLaVA model when using only 80% and 65% of the latency budget, respectively.
Importantly, it consistently adheres to specified latency constraints and generates content-aware execution plans tailored to input images. Furthermore, we show that AdaLLaVA can be integrated with existing token selection techniques designed to enhance efficiency, making it a versatile solution for adaptive inference in MLLMs.
Our key contributions are summarized as follows.
  1. 1.
@@ 2 Related Work
Adaptive inference. Adaptive inference refers to the capability in which the computational complexity of making predictions is dynamically adjusted based on the input data, latency budget, or desired accuracy level [18]. Early works focused on the selection of hand-crafted features in multi-stage prediction pipelines [25, 62, 15]. More recent works have extended these ideas to deep models. For convolutional networks, methods have been developed to selectively downsample inputs, skip layers, or exit early during inference [12, 31, 55, 4, 59, 19, 24, 42, 61, 60]. For vision Transformers, various approaches have been proposed to select different image patches [56, 46, 44], or choose different attention heads and blocks [41, 26]. Similar ideas have also been explored for LLMs and recently MLLMs, where models selectively process tokens [47, 67] or execute a subset of the operations [11, 48] during inference.
Our approach is conceptually similar to existing methods by dynamically selecting a subset of model components during inference. Yet unlike prior methods, our work specifically targets the latency-aware inference of MLLMs, predicting feasible execution plans tailored for input while adhering to varying latency budgets.
@@
While our approach also aims to improve the efficiency of MLLMs, it focuses on dynamically adjusting an MLLM to fit varying latency budgets during inference. This makes our approach orthogonal to prior efforts on developing inherently more efficient MLLMs. Through our experiments, we will demonstrate that our approach is compatible with lightweight models and integrates seamlessly with existing token-pruning techniques (e.g., [49, 8]).
@@ 3.1 Preliminaries: MLLMs
Figure 2: Overview of AdaLLaVA. (a) Model architecture: Our latency encoder embeds an input latency budget into a latency token, which is further processed by the early part of the LLM. The resulting embedding is then fed into the scheduler, leading to the output of an execution plan that controls individual operations in the remaining part of the LLM. Our latency encoder and scheduler are jointly trained with the MLLM. (b) AdaLLaVA-L: the scheduler controls the execution of entire Transformer blocks. (c) AdaLLaVA-H: the scheduler controls the execution of attention heads and MLP neurons, by masking out their activation values and the corresponding weights.
@@ 3.3 Learning to Schedule Execution Plans
Deterministic modeling. One possible solution is to solve a relaxed version of the constrained optimization at training time. We initially explored this solution, where we task $g(\cdot)$ to predict a hard execution plan with binary switches $\mathbf{s}$ and attribute latency violation as part of the objective. This leads to the following loss
@@ 3.5 Model Instantiation
Model architecture. Our goal is to design a lightweight scheduler that minimizes computational overhead yet remains expressive enough to support effective decision-making. To this end, we reuse part of the LLM $f(\cdot)$ to extract visual-language features and encode the latency constraint for the scheduler. Specifically, we first design a latency encoder that converts a latency budget into a token embedding, which is then appended to the original input sequence before being processed by the LLM layers. Within the LLM, the latency token is processed by a few Transformer blocks, attending to all visual-language tokens. The processed token is then passed to a lightweight scheduler that generates the execution plan for the rest of the LLM. Notably, the first few Transformer blocks in the LLM serve two purposes: it simultaneously processes regular MLLM tasks and learns resource allocation based on both content and budget constraints. This design is depicted in Fig. 2 (a).
@@ 4.1 Experimental Setup
Experiment protocol. In most of our experiments, we build on LLaVA-1.5 [36]. Training LLaVA and many other MLLMs typically involves two stages: (1) vision-language alignment pre-training; and (2) visual instruction tuning. We focus on the second stage and seek to jointly finetune the LLM within the MLLM and train our scheduler using visual instruction data, while keeping the vision encoder frozen. Once trained, we perform zero-shot inference across multiple benchmarks following the common practice in the community [36], yet under varying latency budgets.
@@
Baselines and model variants. We mainly compare our model with base model LLaVA-1.5 [36]. We evaluate AdaLLaVA with 7B and 13B (see supplement) models, and across two different designs: (a) AdaLLaVA-L for selecting Transformer blocks; and (b) AdaLLaVA-H for selecting attention heads and MLP activations. In our additional analyses, we also use Mipha-3B [70] as the base model.
@@ 4.2 Main Results
Importantly, for all results, AdaLLaVA adheres to the specified latency budgets (see Sec. 4.3). We provide results on additional VQA benchmarks in Supp. B Tab. A.
Integration with token selection. Token selection techniques have demonstrated recent success in improving the efficiency of MLLMs [49, 8]. AdaLLaVA presents an orthogonal direction in adaptive inference. We now demonstrate that AdaLLaVA can be integrated with token selection to further enhance the efficiency. We combine AdaLLaVA with PruMerge+ [49] and FastV [8], two latest token selection methods designed for MLLMs. For FastV, we set filtering layer KK=2 and filtering ratio RR=50% to ensure consistent comparison. The results are shown in Tab. 1.
With the integration of PruMerge+ or FastV, AdaLLaVA shows significantly improved efficiency across board, when compared to AdaLLaVA with LLaVA-1.5. Again, AdaLLaVA adapts to varying latency budgets and achieves competitive performance relative to the base model (PruMerge+/FastV). For example, with PruMerge+, AdaLLaVA-H shows 2.45% average performance boost at 85% compute budget and only 1.01% performance drop at 60%. A surprising observation is that AdaLLaVA-H achieves strong performance at 85% latency budget, sometimes beating the base model with token pruning. Overall, our results suggest that AdaLLaVA complements to existing token selection approaches. When integrated with Prumerge+ at an 85% latency budget, our approach reduces computational requirements by 70% while maintaining performance with only 1.7% drop in accuracy.
@@ 4.3 Additional Analyses
Our results show that AdaLLaVA can empower a base MLLM with static compute footprint (i.e., LLaVA-1.5, PruMerge+, or FastV as individual dots in the Fig. 3) to adapt to varying accuracy-latency tradeoffs (i.e., the corresponding curves in Fig. 3). With varying latency budgets from 50% to 100%, AdaLLaVA effectively trades compute with accuracy. Integrating with token selection methods (PruMerge+ / FastV) further improves the overall efficiency. Thanks to our sampling process in the probabilistic modeling, AdaLLaVA maintains 0% latency violation. We provide additional visualization of execution plans with different latency in Fig. E in Supp. D.
@@
We visualize attention maps from the latency token to all input visual tokens, computed right before the latency token is fed into the scheduler. This is shown in Fig. 4. These attention maps highlight key regions in the input image for answering the target question. For example, in the top, attention concentrates on “Yes Man” for the movie title question but shifts to the actor name for actor identification question. Further, we visualize the execution plans of different input content given by our scheduler in Fig. 5. Under the same latency budget, AdaLLaVA generates distinct execution plans conditioned on the different visual content. These results show AdaLLaVA’s ability to dynamically adjust its computational focus based on the input image and text query. See our Supp. D for additional visualizations.
Generalization across MLLMs. We further demonstrate that AdaLLaVA can generalize to other MLLMs beyond LLaVA. We consider Mipha-3B [70], a lightweight MLLM built on Phi-2.7B [23]. Specifically, we apply AdaLLaVA-L on Mipha-3B, following its training strategy [70], and report the results on MME benchmark, shown in Tab. 2. These results have similar trend to those with LLaVA-1.5 in Tab. 1. Complete results are presented in Fig. B in Supp. B.
@@
Additional ablations. Ablations on (1) the number and granularity of switches; (2) different designs of switches (i.e., AdaLLaVA-H vs. AdaLLaVA-L); and (3) sampling strategies are included in Supp. C due to space limit.
@@ 5 Conclusion and Discussion
Adaptive inference of MLLMs. Unlike LLMs, MLLMs include a vision encoder and process a large number of redundant visual tokens. While our paper focuses on the scheduling of the LLM component, this adaptivity can be further extended to token selection and vision encoder. We hope this work will be a step toward making MLLMs more viable for real-world applications where computational resources may be constrained and fluctuate significantly.
Relationship to other efficiency methods. This paper explores adaptive inference in MLLMs, emphasizing adaptability to varying latency budgets within a single model. Our approach is orthogonal to prior methods aimed at improving inference efficiency, such as sparse attention [9] and token selection [49]. Indeed, many of these techniques (e.g. token selection as shown in the paper) can be integrated with our framework to further enhance efficiency.
Practical deployment. Our work focuses on algorithm-level innovation, leaving system-level optimization as future work. Conceptually, serving AdaLLaVA is similar to serving MoE-based LLMs [20], which also dynamically routes tokens to different execution paths based on the input. We express compute budgets as percentages of base model FLOPs to abstract hardware/software variations, leaving cross-device portability to future work. We invite joint effort from the vision, learning, and systems communities to further explore these directions.
Acknowledgment. This research was supported in part by the National Science Foundation under Grant Numbers CNS 2333487 / 2333491 (CPS Frontier), CNS 2146449 (CAREER), and IIS 2442739 (CAREER), by the Army Research Lab under contract number W911NF-2020221, and by gift funding from Google and AWS. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@
In this supplementary material, we provide (1) additional implementation details (see Supp. A); (2) detailed results, including broader benchmarks and other LLM backbones on AdaLLaVA accompanying our experiments in Sec. 4 (see Supp. B); (3) further ablations on design of switches and comparison with naive sampling strategies (see Supp. C); (4) additional qualitative results on latency and content adaptivity (see Supp. D) and (5) further discussion on practical deployment (see Supp. E). We hope that this document will complement our main paper.
For sections, figures and equations, we use numbers (e.g., Sec. 1) to refer to the main paper and capital letters (e.g., Sec. A) to refer to this supplement.